OpenStudy (anonymous):

how is principle of mathematical induction actually a prove? lets say we have summation of something = a conjecture what are we actually proving? i dont really see it..

OpenStudy (anonymous):

Good question. Mathematical induction proves the conjecture by first showing it is true for one case, then showing that for any case, the truth of that case implies the truth of the next case.

OpenStudy (anonymous):

ok but there are is a right hand side equals to a left hand side.. how is it actually proving? proving what is true? the left or the right?

OpenStudy (anonymous):

The example you gave: summation of something = a conjecture Is not a proof by induction.

OpenStudy (anonymous):

yes ok..

OpenStudy (anonymous):

It's just a conjecture that you might be able to prove using induction.

OpenStudy (anonymous):

but why does a conjecture have two sides

OpenStudy (anonymous):

what is the summation there for?

OpenStudy (anonymous):

Because you (in this case) are trying to PROVE that the sum of some sequence equals something

OpenStudy (anonymous):

So the equal sign is there because that's what you're assuming is true.

OpenStudy (anonymous):

It could be a less than, or greater than sign.

OpenStudy (anonymous):

it could be all sorts of things.

OpenStudy (anonymous):

so we're trying to prove that the statement itself is true?

OpenStudy (anonymous):

Just depends on what we want to show.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

how did that statement even come about?

OpenStudy (anonymous):

Induction is perfectly reasonable. If \(P(1)\) and \(P(n) \implies P(n+1)\) Then for any \(k\) we can get from \(P(1)\) to \(P(k)\) Suppose we want to show something is true for some \(k\). Since \(P(1)\) and \(P(n) \implies P(n+1)\) then \[P(1) \implies P(2)\]\[P(2)\implies P(3)\]... \[P(n -1)\implies P(n)\]

OpenStudy (anonymous):

A lot of times, it starts off with a guess. We might notice something that seems reasonable. We can make a guess, then prove that it's true with induction.

OpenStudy (anonymous):

ok.. i appreciate your answer but i dont seem to get it in symbol terms seems to confuse me

OpenStudy (anonymous):

P is any proposition you want to show holds for all natural numbers.

OpenStudy (anonymous):

whats propostion

OpenStudy (anonymous):

Something that might be true.

OpenStudy (anonymous):

It is a logical/mathematical statement.

OpenStudy (anonymous):

ok...... i have this question can you go through with me the steps to prove?

OpenStudy (anonymous):

We can certainly try.

OpenStudy (anonymous):

it's summation from r=1 to n r^2= 106n(n+1)(2n+1)

OpenStudy (anonymous):

Sorry correction to the last line of my previous post. Should be: \[P(k-1) \implies P(k)\]

OpenStudy (anonymous):

i mean 1/6

OpenStudy (anonymous):

it's summation from r=1 to n r^2= 1/6n(n+1)(2n+1)

OpenStudy (anonymous):

so what are we showing that is true? the summation part?

OpenStudy (anonymous):

\[\sum_{r=1}^n r^2 = \frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

This?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Ok, so you want to prove it?

OpenStudy (anonymous):

yes! how do i start?

OpenStudy (anonymous):

Start by showing the base case is true.

OpenStudy (anonymous):

Show its true for n=1

myininaya (myininaya):

\[\sum_{i=1}^{1}r^2=1^2=1=\frac{1(2)(3)}{6}=\frac{1(1+1)(2(1)+1)}{6}\]

OpenStudy (anonymous):

what is a base case

OpenStudy (anonymous):

The base case is 1

OpenStudy (anonymous):

You start by showing it's true for some initial n. In this case 1 works fine.

myininaya (myininaya):

now assume it is true for some k integer

OpenStudy (anonymous):

to which side the left or the right

myininaya (myininaya):

both

OpenStudy (anonymous):

Show that if it is true for n, it is true for n+1.

OpenStudy (anonymous):

can i just listen to one person talk?

OpenStudy (anonymous):

Certainly.. Do you have a preference?

OpenStudy (anonymous):

i'd prefer polpak hahah

OpenStudy (anonymous):

='(

OpenStudy (anonymous):

sorry to the rest

OpenStudy (anonymous):

bah. Just cause I'm in green ;p

OpenStudy (anonymous):

Not fair

OpenStudy (anonymous):

no.. u started with the equation!

OpenStudy (anonymous):

lets continue and not waste time!

OpenStudy (anonymous):

It's just sad cause this is one of the more fun problems we get, so the good tutors are champing at the bit ;)

OpenStudy (anonymous):

WAIT CAN I ASK AGAIN WHAT ARE WE PROVING?

OpenStudy (anonymous):

haha..

OpenStudy (anonymous):

We are trying to prove this: \[\sum_{r=1}^n r^2 = \frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

which means we want to prove the right equals to left?

OpenStudy (anonymous):

And for completeness (and simplicity later) I'll modify it slightly: \[p(n) :\sum_{r=1}^n r^2 = \frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

Yes exactly. We need to prove that the left equals the right, for any n you want to choose.

OpenStudy (anonymous):

ok! so we start by n=1?

OpenStudy (anonymous):

yes, we have to start with a base case

OpenStudy (anonymous):

base case=smallest integer?

OpenStudy (anonymous):

So plug 1 into the left side, and into the right side and see that they're equal.

OpenStudy (anonymous):

smallest natural number.

OpenStudy (anonymous):

yes they are equal but you cant assume that since they are true for both sides you cant conclude that pattern is true you have to continue proving by plucking more numbers?

OpenStudy (anonymous):

You can't keep plucking numbers. There might be one that it's not true that you don't pick.

OpenStudy (anonymous):

so instead we use k?

OpenStudy (anonymous):

Instead you have to show that it being true for k means it MUST be true for k+1 and then you have it being true for all numbers.

OpenStudy (anonymous):

so k is any number??

OpenStudy (anonymous):

why not k-1 why k+1

OpenStudy (anonymous):

Because the summation starts at 1 and goes t n, so n has to be bigger than (or equal to) 1 to make sense. So we need to prove for 1 and go up.

OpenStudy (anonymous):

if you wanted to subtract 1 you'd have to start at something big, but there's always a next bigger number to start at.

OpenStudy (anonymous):

ok the next statement says assume Pk is true for some kweird sign Z plus

OpenStudy (anonymous):

meaning?

OpenStudy (anonymous):

hehe.. That wierd sign is \(\in\) and it means 'in'. the \(\mathbb{Z}^+\) means positive integers.

OpenStudy (anonymous):

so! k is one of the positive integer?!

OpenStudy (anonymous):

Yes, we want to show that p(k) is true for any k that is a positive integer.

OpenStudy (anonymous):

ok! cool!

OpenStudy (anonymous):

What is p(k)

OpenStudy (anonymous):

uh huh!

OpenStudy (anonymous):

so now instead of n we write all as k?!

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

we are assuming that that is true?

OpenStudy (anonymous):

Actually if you want we can skip writing p(k). But we'll assume it is true.

OpenStudy (anonymous):

ok! cool so next we got to prove the k+1 term??

OpenStudy (anonymous):

my book says we have to write it..

OpenStudy (anonymous):

ok fair enough.

OpenStudy (anonymous):

lets prove the k+1 term now

OpenStudy (anonymous):

Okey

OpenStudy (anonymous):

summation of r=1 to (k+1)???

OpenStudy (anonymous):

why the last term (k+1)

OpenStudy (anonymous):

and why is it want to show

OpenStudy (anonymous):

want to show that is it actually also equals to that??

OpenStudy (anonymous):

Right. We can show that IF we assume it's true for p(k) then it must be true for p(k+1)

OpenStudy (anonymous):

oh you assumed that Pk is true!

OpenStudy (anonymous):

Now we are gonna make the assumption that it's true for p(k) which is the sum from 1 to k

OpenStudy (anonymous):

why = 1+2+3+............